how do you solve this linear system by using substitution?
0.5x+y=9
1.6x+0.2y=13

Answer by stanbon(75887)

0.5x+y=9
1.6x+0.2y=13
----
y = 9-0.5x
---
Substitute for "y" and solve for "x":
1.6x + 0.2(9-0.5x) = 13
1.6x + 1.8 -0.01x = 13
------
1.59x = 11.2
x = 7.044
------
Solve for "y":
y = 9-0.5(7.044)
-----
y = 5.48
------------------
Cheers,
Stan H.
==================